# Borsuk-Ulam and Replicable Learning of Large-Margin Halfspaces

### v3

@@ 3.2 The Upper Bound
Let $\bm{w}_{1},\dots,\bm{w}_{k}$ be as in Algorithm 1. Since $\mathcal{D}$ is realizable by $\mathcal{H}_{\gamma}^{d}$, for every $i$, $\bm{w}_{i}$ is a $\gamma$-separator for $\mathbf{S}_{i}$. Therefore, by Theorem 3.4, if $n_{0}(\epsilon,\delta,k)$ is sufficiently large,
$\Pr_{\begin{subarray}{c}\mathbf{S}_{i}\sim\mathcal{D}^{n_{0}}\end{subarray}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\right]\geq 1-\frac{\delta}{4k}.$
@@
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\text{ for all $i\in[k]$}\right]\geq 1-\frac{\delta}{4},$ (6)
@@
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\min_{i\in[k]}\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\epsilon\right]\geq 1-\frac{\delta}{4}.$ (7)
Finally, if $(x,y)\in\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$ satisfies $y{\left\langle x,\bm{w}_{i}\right\rangle}\geq\gamma/2$ for all $i\in[k]$, then noting that $\|\bm{w}\|\leq 1$, we have
@@
Thus, from Equation 7, we have
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}\|_{2}<\frac{\gamma}{2}\right]\leq\frac{\delta}{4}$
@@
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}\left[\bm{y}{\left\langle\bm{x},\bm{z}\right\rangle}\geq\frac{\gamma}{2}\right]\geq 1-\epsilon\right]\geq 1-\frac{\delta}{4}.$ (9)
By applying Lemma 3.7 with $\alpha\coloneqq\gamma/2$, after rounding $\bm{z}$ to $\tilde{\bm{z}}$, we have $\|\tilde{\bm{z}}-\bm{z}\|_{2}<\frac{\gamma}{2}$. Thus if $(x,y)\in\mathbb{S}^{d-1}\times\{\pm 1\}$ satisfy $y{\left\langle x,\bm{z}\right\rangle}\geq\gamma/2$, then
$y{\left\langle x,\tilde{\bm{z}}\right\rangle}=y{\left\langle x,\bm{z}\right\rangle}+y{\left\langle x,\tilde{\bm{z}}-\bm{z}\right\rangle}\geq\frac{\gamma}{2}-\|\tilde{\bm{z}}-\bm{z}\|_{2}>0,$
namely $\overline{h}_{\tilde{\bm{z}}}(x)=y$. Thus,
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\mathcal{L}_{\mathcal{D}}(\overline{h}_{\tilde{\bm{z}}})\leq\epsilon\right]\geq 1-\frac{\delta}{4},$
@@
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}-\operatorname{\mathbb{E}}[\bm{w}]\|_{2}\geq\frac{\gamma\beta}{2}\right]\leq\frac{\delta}{4}.$
Since $\bm{z}=\frac{\bm{w}}{\|\bm{w}\|_{2}}$, by applying the union bound to Eq. 4 and the above inequality, we have
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{z}-\operatorname{\mathbb{E}}[\bm{z}]\|_{2}\geq\beta\right]\leq\frac{\delta}{2}.$
Consequently, by Lemma 3.7, with probability at least $21-\delta/$ 2, the rounding scheme ($\mathrm{round}_{\frac{\gamma}{2}}$) outputs one of at most $d$ hypotheses. Applying a union bound with 3.8 completes the proof of the upper bound of Theorem 1.6. ∎
@@ 4 Disambiguations of Gap Hamming Distance
We first prove that every disambiguation of the Gap Hamming Distance matrix family has unbounded public-coin randomized communication complexity.
Proof of Corollary 1.15.
We combine Theorem 1.14 with the known relationship between Littlsetone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity. Given a matrix $M\in\{\pm 1\}^{\mathcal{X}\times\mathcal{Y}}$, the margin of $M$ is defined
@@
Let $\{M_{n}\}_{n=1}^{\infty}$ be a family of disambiguation of $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$. By Theorem 1.14, we know that $\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n})$. It thus follows from LABEL:{Ldim_of_margin} that
@@
$\Omega(\log(\operatorname{m}(M_{n})^{-1}))=\Omega(\log\log n).\qed$
We next present the proof of Theorem 1.14. The key is to use an embedding of bounded margin half-spaces in dimension $d$ into the Boolean cube, which allows us to disambiguate $\mathcal{H}^{d}_{\gamma}$ using a disambiguation of the Gap Hamming Distance problem in dimension $O(d)$. The existence of such an embedding was proved in [HHM23], which we rephrase as follows.
Lemma 4.1 (Adapted from [HHM23, Lemma 3.2]).
$\langle u,v\rangle\geq\gamma^{\prime}\implies\langle\xi(u),\xi(v)\rangle\geq\gamma n$
$\langle u,v\rangle\leq-\gamma^{\prime}\implies\langle\xi(u),\xi(v)\rangle\leq-\gamma n$
We are now ready to prove Theorem 1.14.
Proof of Theorem 1.14.
We will use this lemma along with the functions $\left\{M_{n}\right\}_{n=1}^{\infty}$ to disambiguate the family of partial concept classes $\{\mathcal{H}_{\gamma^{\prime}}^{d(n)}\}_{n=1}^{\infty}$. To this end, we disambiguate each partial concept $h_{w}\in\mathcal{H}_{\gamma^{\prime}}^{d}$ (defined in LABEL:{eq:def_large_margin}) to
$\overline{h}_{w}(x)\coloneqq M_{n}\big{(}\xi(w),\xi(x)\big{)}.$
Let us verify that $\overline{h}_{w}$ is, in fact, a disambiguation of $h_{w}$.
Suppose that $h_{w}(x)=1$. By definition, this occurs exactly when $\langle w,x\rangle\geq\gamma^{\prime}$. It follows from the properties of $\xi$ that
$\langle\xi(w),\xi(x)\rangle\geq\gamma n.$
Therefore, for such $w,x$, we have
$\overline{h}_{w}(x)=M_{n}(\xi(w),\xi(x))=\operatorname{GHD}_{\gamma}^{n}(\xi(w),\xi(x))=1=h_{w}(x).$
A similar argument shows that if $h_{w}(x)=-1$, then $\overline{h}_{w}(x)=-1$. We deduce that $\overline{h}_{w}$ indeed disambiguates $h_{w}$, and ¯$\overline{\mathcal{H}}^{d}_{\gamma^{\prime}}$ is a disambiguation of $\mathcal{H}^{d}_{\gamma^{\prime}}$.
Finally, note that by construction, any shattered mistake tree in ¯$\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}$ corresponds to a shattered mistake tree of the same depth in $M_{n}$. Therefore, $\operatorname{Ldim}(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d})\leq\operatorname{Ldim}(M_{n})$. This combined with Theorem 1.8 implies that
$\operatorname{Ldim}(M_{n})\geq\operatorname{Ldim}\left(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}\right)=\Omega\left(\sqrt{\log d(n)}\right)=\Omega\left(\sqrt{\log n}\right).\qed$
Appendix A Equivalence of Stability and List-Replicability
We prove the following more refined version of Theorem 1.2.
Theorem A.1.
@@ Appendix A Equivalence of Stability and List-Replicability
Let $\mathcal{H}$ be any total or partial concept class on the domain $\mathcal{X}$. Then for every $\epsilon\in(0,1)$,
$\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}\qquad\text{ and }\qquad\operatorname{LR}_{\epsilon}(\mathcal{H})\leq\frac{1}{\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{H})}.$
Consequently, $\rho^{\mathrm{gs}}(\mathcal{H})=\frac{1}{\operatorname{LR}(\mathcal{H})}$.
@@
We first prove that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}$. Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon,L)$-list replicable learner for $\mathcal{H}$ with sample complexity $n=n(\epsilon,\delta)$. Let $\mathcal{D}$ be any realizable distribution on $\mathcal{X}\times\{\pm 1\}$, and let $h_{1},\ldots,h_{L}$ be the list of hypotheses guaranteed by Definition 1.1.
@@
$\Pr_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\bm{S})=h_{i}]\geq\frac{1-\delta}{L}.$
Since this statement holds for arbitrary $\delta>0$, $\bm{\mathcal{A}}$ is itself an $(\epsilon,\rho)$-globally stable learner for all $\rho<\frac{1}{L}$. We may conclude that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{H})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{H})}$.
Next, we prove that $\operatorname{LR}_{\epsilon}(\mathcal{H})\leq 1/\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{H})$. Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon/3,\rho)$-globally stable learner for $\mathcal{H}$ with sample complexity $n_{0}=n_{0}(\epsilon)$. By the stability assumption, for every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\{\pm 1\}$, there exists $h^{*}:\mathcal{X}\to\left\{\pm 1\right\}$ satisfying
$\mathcal{L}_{\mathcal{D}}(h^{*})\leq\frac{\epsilon}{3}\ \text{ and }\ \Pr_{\begin{subarray}{c}\bm{S}\sim\mathcal{D}^{n_{0}}\end{subarray}}[\bm{\mathcal{A}}(\bm{S})=h^{*}]\geq\rho.$ (10)
For every $h\in\{\pm 1\}^{\mathcal{X}}$ and realizable distribution $\mathcal{D}$, define
$p(h)\coloneqq\Pr_{\mathbf{S}\sim\mathcal{D}^{n_{0}}}[\bm{\mathcal{A}}(\bm{S})=h],$
@@
$\Lambda\coloneqq\left\{h\in\left\{\pm 1\right\}^{\mathcal{X}}\ |\ p(h)>\frac{1}{L+1}\text{ and }\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right\}.$
@@
$\bm{S}=({\bm{P}},\bm{Q})\sim\mathcal{D}^{tn_{0}+n_{1}},\quad\text{where }{\bm{P}}=({\bm{P}}_{1},\ldots,{\bm{P}}_{t})\sim(\mathcal{D}^{n_{0}})^{t}=\mathcal{D}^{tn_{0}},\quad\text{and}\quad\bm{Q}\sim\mathcal{D}^{n_{1}}.$
@@
  • $\mathcal{L}_{{\bm{Q}}}(h)\leq\tfrac{2\epsilon}{3}$
@@
$A:\quad$ $|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}$
@@
$\Pr_{{\bm{P}}\sim\mathcal{D}^{tn_{0}}}[B]\geq 1-\frac{\delta}{2}\quad\text{and}\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}[A]\geq 1-\frac{\delta}{2}.$
Proof of A.2.
For the choice of $t$ and the proof of the first inequality, we use the uniform convergence property of the family of indicator functions on {±$\{\pm 1\}^{\mathcal{X}}$. More precisely, for $f\in\left\{\pm 1\right\}^{\mathcal{X}}$, define $\mathbb{I}_{f}:\left\{\pm 1\right\}^{\mathcal{X}}\to\left\{0,1\right\}$ as
@@
has VC dimension $1$, and therefore, it satisfies the uniform convergence property. For ${\bm{P}}_{i}\sim\mathcal{D}^{n_{0}}$, $\bm{\mathcal{A}}({\bm{P}}_{i})$ induces a probability distribution $\mu$ on {±$\left\{\pm 1\right\}^{\mathcal{X}}$, and we have
$1-p(h)=\Pr_{{\bm{P}}_{i}\sim\mathcal{D}^{n_{0}}}[\bm{\mathcal{A}}({\bm{P}}_{i})\neq h]=\mathcal{L}_{\mu}(\mathbb{I}_{h}),$
@@
Now that we have $t$, we can define $n_{1}$ and prove the second inequality. Note that for every $h\in\{\pm 1\}^{\mathcal{X}}$, for ${\bm{Q}}\sim\mathcal{D}^{n_{1}}$, $\mathcal{L}_{\bm{Q}}(h)$ is an average of ${n_{1}}$ samplings of a Bernoulli random variable with expectation $\mathcal{L}_{\mathcal{D}}(h)$. Thus, by Hoeffding’s inequality, there exists $n_{1}=n_{1}(\epsilon^{\prime},t)$ such that
$\Pr_{{\bm{Q}}\sim\mathcal{D}^{n_{1}}}\left[\left|\mathcal{L}_{\mathcal{D}}(h)-\mathcal{L}_{\bm{Q}}(h)\right|>\frac{\epsilon}{3}\right]\leq\frac{\delta}{2t}.$ (11)

### v4

@@ 3.2 The upper bound
Let $\bm{w}_{1},\dots,\bm{w}_{k}$ be as in Algorithm 1. Since $\mu$ is realizable by $\mathcal{H}_{\gamma}^{d}$, for every $i$, $\bm{w}_{i}$ is a $\gamma$-separator for $\mathbf{S}_{i}$. Therefore, by Theorem 3.4, if $n_{0}(\epsilon,\delta,k)$ is sufficiently large,
$\Pr_{\begin{subarray}{c}\mathbf{S}_{i}\sim\mu^{n_{0}}\end{subarray}}\left[\Pr_{(\bm{x},\bm{y})\sim\mu}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\right]\geq 1-\frac{\delta}{4k}.$
@@
$\Pr_{\mathbf{S}\sim\mu^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mu}\left[\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\frac{\epsilon}{k}\text{ for all $i\in[k]$}\right]\geq 1-\frac{\delta}{4},$ (7)
@@
$\Pr_{\mathbf{S}\sim\mu^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mu}\left[\min_{i\in[k]}\bm{y}{\left\langle\bm{x},\bm{w}_{i}\right\rangle}<\frac{\gamma}{2}\right]\leq\epsilon\right]\geq 1-\frac{\delta}{4}.$ (8)
[…]
@@
Thus, from (8), we have
[…]
@@
$\Pr_{\mathbf{S}\sim\mu^{kn_{0}}}\left[\Pr_{(\bm{x},\bm{y})\sim\mu}\left[\bm{y}{\left\langle\bm{x},\bm{z}\right\rangle}\geq\frac{\gamma}{2}\right]\geq 1-\epsilon\right]\geq 1-\frac{\delta}{4}.$ (10)
[…]
namely $\overline{c}_{\tilde{\bm{z}}}(x)=y$. Thus,
$\Pr_{\mathbf{S}\sim\mu^{kn_{0}}}\left[\operatorname{loss}_{\mu}(\overline{c}_{\tilde{\bm{z}}})\leq\epsilon\right]\geq 1-\frac{\delta}{4},$
@@
[…]
Consequently, by Lemma 3.7, with probability at least $1-\delta/2$, the rounding scheme ($\mathrm{round}_{\frac{\gamma}{2}}$) outputs one of at most $d$ hypotheses. Applying a union bound with 3.8 completes the proof of the upper bound of Theorem 2.1. ∎
@@ 4 Disambiguations of gap Hamming distance
We prove Theorem 2.5 in this section.
The Littlestone dimension of disambiguations.
The key to obtaining (3) is to use an embedding of bounded margin half-spaces in dimension $d$ into the Boolean cube, which allows us to disambiguate $\mathcal{H}^{d}_{\gamma}$ using a disambiguation of the Gap Hamming Distance problem in dimension $O(d)$. The existence of such an embedding was proved in [HHM23], which we rephrase as follows.
Lemma 4.1 (Adapted from [HHM23, Lemma 3.2]).
${\left\langle u,v\right\rangle}\geq\gamma^{\prime}\implies{\left\langle\xi(u),\xi(v)\right\rangle}\geq\gamma n$
${\left\langle u,v\right\rangle}\leq-\gamma^{\prime}\implies{\left\langle\xi(u),\xi(v)\right\rangle}\leq-\gamma n$
We will use this lemma along with the functions $\left\{M_{n}\right\}_{n=1}^{\infty}$ to disambiguate the family of partial concept classes $\{\mathcal{H}_{\gamma^{\prime}}^{d(n)}\}_{n=1}^{\infty}$. To this end, we disambiguate each partial concept $c_{w}\in\mathcal{H}_{\gamma^{\prime}}^{d}$ (defined in (1)) to
$\overline{c}_{w}(x)\coloneqq M_{n}\big(\xi(w),\xi(x)\big).$
Let us verify that $\overline{c}_{w}$ is, in fact, a disambiguation of $c_{w}$.
Suppose that $c_{w}(x)=1$. By definition, this occurs exactly when ${\left\langle w,x\right\rangle}\geq\gamma^{\prime}$. It follows from the properties of $\xi$ that
${\left\langle\xi(w),\xi(x)\right\rangle}\geq\gamma n.$
Therefore, for such $w,x$, we have
$\overline{c}_{w}(x)=M_{n}(\xi(w),\xi(x))=\operatorname{GHD}_{\gamma}^{n}(\xi(w),\xi(x))=1=c_{w}(x).$
A similar argument shows that if $c_{w}(x)=-1$, then $\overline{c}_{w}(x)=-1$. We deduce that $\overline{c}_{w}$ indeed disambiguates $c_{w}$, and $\overline{\mathcal{H}}^{d}_{\gamma^{\prime}}$ is a disambiguation of $\mathcal{H}^{d}_{\gamma^{\prime}}$.
Finally, note that by construction, any shattered mistake tree in $\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}$ corresponds to a shattered mistake tree of the same depth in $M_{n}$. Therefore, $\operatorname{Ldim}(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d})\leq\operatorname{Ldim}(M_{n})$. This combined with Theorem 2.3 implies that
$\operatorname{Ldim}(M_{n})\geq\operatorname{Ldim}\left(\overline{\mathcal{H}}_{\gamma^{\prime}}^{d}\right)=\Omega\left(\sqrt{\log d(n)}\right)=\Omega\left(\sqrt{\log n}\right).$
Communication complexity of disambiguations.
To complete the proof of Theorem 2.5 we use the known relationships between Littlestone dimension, margin, distributional discrepancy, and public-coin randomized communication complexity. Given a matrix $M\in\{\pm 1\}^{\mathcal{X}\times\mathcal{Y}}$, the margin of $M$ is defined
@@
Let $\{M_{n}\}_{n=1}^{\infty}$ be a family of disambiguation of $\{\mathrm{GHD}_{\gamma}^{n}\}_{n=1}^{\infty}$. By (3), we know that $\operatorname{Ldim}(M_{n})=\Omega(\sqrt{\log n})$. It thus follows from (2) that
@@
$\Omega(\log(\operatorname{m}(M_{n})^{-1}))=\Omega(\log\log n).$
Acknowledgments
We are grateful to Arkadev Chattopadhyay for pointing out the connection to pseudo-determinism and for valuable comments on the exposition. We also thank Zachary Chase for clarifying the implicit bound in [GGKM21].
Appendix A Replicability and privacy notions
In this section, we state the formal definitions of shared-randomness replicability, DP-learnability, and global stability.
A.1 Shared-randomness replicability
Let $\mathcal{A}(S,r)$ be a randomized learning rule, where $r$ denotes the random seed.
Definition A.1 (Shared-randomness replicability [GKM21, ILPS22]).
A concept class $\mathcal{C}\subseteq\left\{\pm,\star\right\}^{\mathcal{X}}$ is shared-randomness replicable if there exists a learning rule $\mathcal{A}$ and a sample complexity function $n(\epsilon,\delta)$ such that, for every $\epsilon,\delta>0$ and every realizable distribution $\mu$, the following conditions hold:
  • Small population loss: $\Pr_{\bm{S}\sim\mu^{n},\bm{r}}[\operatorname{loss}_{\mu}\left(\mathcal{A}(\bm{S},\bm{r})\right)>\epsilon]\leq\delta.$
  • Replicability with shared randomness: $\Pr_{\bm{S,S^{\prime}}\sim\mu^{n},\bm{r}}[\mathcal{A}(\bm{S},\bm{r})=\mathcal{A}(\bm{S^{\prime}},\bm{r})]\geq 1-\delta.$
One could consider shared-randomness replicability to be a weak form of replicability, as different executions of the algorithm can use the same random seed.
A.2 Differential privacy
[…]
Definition A.2 (Differential privacy).
Given $\epsilon,\delta>0$, a randomized learning rule
In approximate differential privacy, the parameters $\epsilon$ and $\delta$ are typically set as follows: $\epsilon$ is taken to be a small fixed constant (e.g., 0.1), while $\delta$ is a negligible function, $\delta=n^{-\omega(1)}$.
Definition A.3 (Approximate differentially drivate learnability).
We say that a concept class $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is approximate differentially private learnable (DP-learnable) if there is a learning rule $\bm{\mathcal{A}}:(\mathcal{X}\times\left\{\pm 1\right\})^{*}\to\left\{\pm 1\right\}^{\mathcal{X}}$ with sample complexity $n(\epsilon,\delta)$ such that for every $\epsilon,\delta>0$ the following holds.
  • The class $\mathcal{C}$ is $(\epsilon,\delta)$-PAC learnable by $\bm{\mathcal{A}}$ using $n(\epsilon,\delta)$ samples.
  • The learning rule $\bm{\mathcal{A}}$ applied to samples of size $n(\epsilon,\delta)$ is $(\epsilon^{\prime}(n),\delta^{\prime}(n))$-differentially private learnable where $\epsilon^{\prime}(n)\leq 0.1$ and $\delta^{\prime}(n)\leq n^{-w(1)}$.
Definition A.4 (Pure Differentially Private Learnability).
We say that a concept class $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is pure differentially private learnable (pure DP-learnable) if $\mathcal{C}$ is PAC learnable by a $(0.1,0)$-differentially private learning rule.
A.3 Global stability
Definition A.5.
A learning rule $\bm{\mathcal{A}}$ for a concept class $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is $(\epsilon,\rho)$-globally stable if for every realizable distribution $\mu$, there is a hypothesis $h\in\left\{\pm 1\right\}^{\mathcal{X}}$ with population loss $\operatorname{loss}_{\mu}(h)\leq\epsilon$ satisfying
$\Pr_{\bm{S}\sim\mu^{n}}[\bm{\mathcal{A}}(\bm{S})=h]\geq\rho,\ \ \text{ where }n=n(\epsilon).$
We define $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{C})$ to be the supremum of $\rho$ such that there is a $(\epsilon,\rho)$-globally stable learner for $\mathcal{C}$. The global stability parameter of $\mathcal{C}$ is then defined as
$\rho^{\mathrm{gs}}(\mathcal{C})\coloneqq\inf_{\epsilon>0}\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{C}).$
The definition of global stability might initially seem weak, as a globally stable learner is not necessarily a PAC learner. In particular, since $\rho$ can be a small constant, there may be a probability as great as $1-\rho$ that the learning rule outputs a hypothesis with large population loss. However, as discussed in the next section, global stability is equivalent to the seemingly stronger notion of list replicability.
Appendix B Equivalence of global stability and list replicability
In [CMY23], Chase, Moran and Yehudayoff proved that for every total class $\mathcal{C}\subseteq\left\{\pm 1\right\}^{\mathcal{X}}$, list replicability is equivalent to global stability. It is easy to check that their proof applies to partial concept classes, resulting in the following relationship between the list replicability number and the global stability parameter.
Theorem B.1.
@@ Appendix B Equivalence of global stability and list replicability
Let $\mathcal{C}$ be any total or partial concept class on the domain $\mathcal{X}$. Then for every $\epsilon\in(0,1)$,
$\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{C})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{C})}\qquad\text{ and }\qquad\operatorname{LR}_{\epsilon}(\mathcal{C})\leq\frac{1}{\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{C})}.$
Consequently, $\rho^{\mathrm{gs}}(\mathcal{C})=\frac{1}{\operatorname{LR}(\mathcal{C})}$.
@@
We first prove that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{C})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{C})}$. Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon,L)$-list replicable learner for $\mathcal{C}$ with sample complexity $n=n(\epsilon,\delta)$. Let $\mu$ be any realizable distribution on $\mathcal{X}\times\{\pm 1\}$, and let $h_{1},\ldots,h_{L}$ be the list of hypotheses guaranteed by Definition 1.2.
@@
$\Pr_{\mathbf{S}\sim\mu^{n}}[\bm{\mathcal{A}}(\bm{S})=h_{i}]\geq\frac{1-\delta}{L}.$
Since this statement holds for arbitrary $\delta>0$, $\bm{\mathcal{A}}$ is itself an $(\epsilon,\rho)$-globally stable learner for all $\rho<\frac{1}{L}$. We may conclude that $\rho^{\mathrm{gs}}_{\epsilon}(\mathcal{C})\geq\frac{1}{\operatorname{LR}_{\epsilon}(\mathcal{C})}$.
Next, we prove that $\operatorname{LR}_{\epsilon}(\mathcal{C})\leq 1/\rho^{\mathrm{gs}}_{\epsilon/3}(\mathcal{C})$. Let $\epsilon>0$ be an accuracy parameter, and let $\bm{\mathcal{A}}$ be an $(\epsilon/3,\rho)$-globally stable learner for $\mathcal{C}$ with sample complexity $n_{0}=n_{0}(\epsilon)$. By the stability assumption, for every realizable distribution $\mu$ on $\mathcal{X}\times\{\pm 1\}$, there exists $h^{*}:\mathcal{X}\to\left\{\pm 1\right\}$ satisfying
$\operatorname{loss}_{\mu}(h^{*})\leq\frac{\epsilon}{3}\ \text{ and }\ \Pr_{\begin{subarray}{c}\bm{S}\sim\mu^{n_{0}}\end{subarray}}[\bm{\mathcal{A}}(\bm{S})=h^{*}]\geq\rho.$ (11)
For every $h\in\{\pm 1\}^{\mathcal{X}}$ and realizable distribution $\mu$, define
$p(h)\coloneqq\Pr_{\mathbf{S}\sim\mu^{n_{0}}}[\bm{\mathcal{A}}(\bm{S})=h],$
@@
$\Lambda\coloneqq\left\{h\in\left\{\pm 1\right\}^{\mathcal{X}}\ |\ p(h)>\frac{1}{L+1}\text{ and }\operatorname{loss}_{\mu}(h)\leq\epsilon\right\}.$
@@
$\bm{S}=({\bm{P}},\bm{Q})\sim\mu^{tn_{0}+n_{1}},\quad\text{where }{\bm{P}}=({\bm{P}}_{1},\ldots,{\bm{P}}_{t})\sim(\mu^{n_{0}})^{t}=\mu^{tn_{0}},\quad\text{and}\quad\bm{Q}\sim\mu^{n_{1}}.$
@@
  • $\operatorname{loss}_{{\bm{Q}}}(h)\leq\tfrac{2\epsilon}{3}$
@@
$A:\quad$ $|\operatorname{loss}_{\mu}(h)-\operatorname{loss}_{\bm{Q}}(h)|\leq\frac{\epsilon}{3}\text{ for all }h\in\mathcal{Y}$
@@
$\Pr_{{\bm{P}}\sim\mathcal{D}^{tn_{0}}}[B]\geq 1-\frac{\delta}{2}\quad\text{and}\Pr_{{\bm{Q}}\sim\mu^{n_{1}}}[A]\geq 1-\frac{\delta}{2}.$
Proof of B.2.
For the choice of $t$ and the proof of the first inequality, we use the uniform convergence property of the family of indicator functions on $\{\pm 1\}^{\mathcal{X}}$. More precisely, for $f\in\left\{\pm 1\right\}^{\mathcal{X}}$, define $\mathbb{I}_{f}:\left\{\pm 1\right\}^{\mathcal{X}}\to\left\{0,1\right\}$ as
@@
has VC dimension $1$, and therefore, it satisfies the uniform convergence property. For ${\bm{P}}_{i}\sim\mu^{n_{0}}$, $\bm{\mathcal{A}}({\bm{P}}_{i})$ induces a probability distribution $\mu$ on $\left\{\pm 1\right\}^{\mathcal{X}}$, and we have
$1-p(h)=\Pr_{{\bm{P}}_{i}\sim\mu^{n_{0}}}[\bm{\mathcal{A}}({\bm{P}}_{i})\neq h]=\operatorname{loss}_{\mu}(\mathbb{I}_{h}),$
@@
Now that we have $t$, we can define $n_{1}$ and prove the second inequality. Note that for every $h\in\{\pm 1\}^{\mathcal{X}}$, for ${\bm{Q}}\sim\mu^{n_{1}}$, $\operatorname{loss}_{\bm{Q}}(h)$ is an average of $n_{1}$} samplings of a Bernoulli random variable with expectation $\operatorname{loss}_{\mu}(h)$. Thus, by Hoeffding’s inequality, there exists $n_{1}=n_{1}(\epsilon^{\prime},t)$ such that
$\Pr_{{\bm{Q}}\sim\mu^{n_{1}}}\left[\left|\operatorname{loss}_{\mu}(h)-\operatorname{loss}_{\bm{Q}}(h)\right|>\frac{\epsilon}{3}\right]\leq\frac{\delta}{2t}.$ (12)